# Starlight-driven flared-staircase geometry in radiation hydrodynamic models of protoplanetary disks

### v1

@@ Abstract
Context. Protoplanetary disks observed in millimeter continuum and scattered light show a variety of substructures. While embedded planets are a common explanation, during the early planet formation phase, various physical processes in the disk could also trigger such features. One such possibility that has been previously theorized for passive disks is the the thermal wave instability or the stellar irradiation instability—the flared disk may become unstable as directly illuminated regions puff up and cast shadows behind them. This would manifest as bright and dark rings, and a staircase-like structure in the disk optical surface.
@@
Conclusions. We conclude that we can recover thermally induced flared-staircase structure in radiation hydrodynamic simulations of irradiated protoplanetary disks using the flux-limited diffusion method. We find that the shadowing effect is sensitive to the dust content in the disk. We highlight the importance of modeling dust dynamics consistently to explain starlight-driven shadows.
@@ 1 Introduction
Disk substructures are ubiquitous in modern observations of planet-forming disks across different wavelengths. These include millimeter thermal emission from ground-based instruments like ALMA (2018Andrews), scattered light observations with VLT-SPHERE (2018Avenhaus; 2023Benisty), and more recently, studies with JWST (2025Tazaki). Planets in the disk could carve out gaps (2018Zhang) that explain substructure, of which the planetary gaps in the PDS 70 and WISPIT 2 systems are direct observational constraints (2018Keppler; 2019Haffert; 2025VanCap; 2025Close). Many physical processes in the disk could also trigger substructures (2023Bae), such as zonal flows due to the magneto-rotational instability (MRI, 2011Uribe; 2015Flock), spirals launched by gravitational instability (GI, 2016Kratter), gaps and vortices resulting from the vertical shear instability (VSI, 2020Flock), and spirals and rings as a consequence of misaligned inner disks (2024Zhang; 2025bZiampras) among others. Mechanisms such as these which could invert the local pressure gradient, radially pile up dust in the disk, acting as dust traps (1972Whipple; 2004Paardekooper; 2024Ormel).
@@
The mechanism behind this starlight-driven shadowing is outlined as follows (see also Fig. 1 in 2021Ueda, for an illustration): in vertical hydrostatic equilibrium, the balance between gravity and pressure— set by stellar irradiation heating– yields a passive (irradiated) disk with flared geometry. Thus, outer regions absorb substantially more stellar light than in an internally-heated, conical disk. For optically thick disks, scale height perturbations could trigger thermal waves that travel inwards, puffing up regions of the disk and casting shadows behind them. This shows up as bright and dark bands or shadows in temperature, or staircases in the optical surface. One can refer to them as ”unstable” shadows to distinguish them from the ”stable” shadow cast by the hot inner rim in transition disks (2001Dullemond) or due to an opacity transition at the snow line (2015Bailli). For clarity throughout this article, we adopt the following definitions for the words ”staircase”, ”bump”, and ”shadow”: a staircase is a step-like modulation in the irradiation absorption surface as a function of radius, consisting of a heated, puffed-up front (a bump) and a colder backside (a shadow). In this context, we note that a bump refers to the locally elevated, irradiated region, rather than the more traditional temperature/pressure gradient inversion commonly associated with the term. A shadow, in turn, corresponds to the dip in the irradiation flux downstream, which we trace through the midplane temperature decrease of at least 5% relative to the background power-law, this is clearly illustrated by the sketch in Fig. 1.
@@
Figure 1: Schematic to illustrate the different terms: ”bump”, ”shadow” and ”staircase” defined in the introduction. The star is denoted by the orange blob, and the stellar rays hit the irradiation absorption surface $\tau_{\star}=1$, at height $H_{s}=H\penalty 10000\ (\tau_{\star}=1)$ from the midplane, and a flaring angle $\varphi$.
@@
There have been some efforts in modeling the thermal wave instability with 2.5D radiation hydrodynamic simulations in recent literature. 2022Pavlyuchenkov2D used a diffusion solver for radiation transport and 2022FuksmanKlahr used M1 moment transfer coupled with frequency-dependent irradiation. While these works reported inwardly propagating thermal waves and shadows in the hydrostatic models, both studies concluded that the instability is damped in dynamical simulations. 2022FuksmanKlahr attributed this to the inefficiency in transporting thermal perturbations from the atmosphere to the midplane in the disk, as radial advection is more dominant. 2024Kutra used a more simple ”radiative forcing” method which enforced the locally isothermal approximation and found staircase structures that did not propagate inwards but settled in the disk for several orbital timescales.
@@
In Sects. 2 & 2.4 we outline the physics used in our model, the model parameters, and the numerical methods used. In Sects. 35 we show the results from our hydrostatic, dynamical, and dust models. We then discuss the implications of our results in the context of previous work in Sect. 6 and summarize our findings in Sect. 7.
@@ 2.1 Radiation hydrodynamic equations
The gravitational potential of the star is given by $\Phi_{\star}=-\text{G}M_{\star}/r$ where G is the gravitational constant, $M_{\star}$ is the mass of the central star, and $r=\sqrt{R^{2}+z^{2}}$ is the spherical radius distance to the star. The total energy is a function of kinetic energy and the specific internal energy $e$ with $E_{\text{tot}}=\rho e+\frac{1}{2}\rho\bm{u}^{2}$. The gas pressure is denoted by $P$, and $\bm{\mathrm{I}}$ is the unit matrix. The temperature-dependent Planck and Rosseland mean opacities are $\kappa_{\text{P}}$ and $\kappa_{\text{R}}$ respectively, $a_{\text{R}}$ is the radiation constant, and $c$ is the speed of light. The set of equations is complete with a closure relation, that is, the ideal gas equation $P=(\gamma-1)\rho e$ with the adiabatic index chosen as $\gamma=1.42$ (1978Decampli) and the mean molecular weight $\mu=2.35$. The gas temperature $T$ is given by $e=\rho c_{\text{V}}T$, with $c_{\text{V}}=R_{\text{gas}}/\mu(\gamma-1)$ being the specific heat capacity at constant volume, and $R_{\text{gas}}$ is the gas constant. The isothermal sound speed is defined as $c_{\mathrm{s}}=\sqrt{P/\rho}$ and the gas pressure scale height is given by $H=c_{\mathrm{s}}/\Omega_{\mathrm{K}}$ where $\Omega_{\mathrm{K}}=\sqrt{GM_{\star}/R^{3}}$ is the Keplerian angular velocity.
@@ 2.3 Model parameters
ML1* $0.01$ 11 $4000$ $2.086$ $200$
Table 1: Parameters for the different hydrostatic+dynamic runs. Each model is denoted by M followed by a label indicating the parameter altered relative to the fiducial run MFID. ML1* is an additional model we set up that matches 2022FuksmanKlahr with a stellar parameters corresponding to $L_{\star}=1\,\mathrm{L}_{\odot}$, that we only discuss during the comparison in Appendix A.
@@ 4 Results II: Changing the dust content
In this section, we outline our results from our hydrostatic and dynamical models for the low dust-to-gas ratio case of $\epsilon_{\text{dg}}=0.001$. As outlined above, these values are more typical to the dust content observed in later stages (i.e., Class II protoplanetary disks 2021Rilinger; 2024Birnstiel). Figure 7 shows the temperature and the irradiation source term after 30 iterations for the hydrostatic model (top) and the hydrodynamic model after $t=500$ orbits. We do not see any temperature bumps/shadows forming in the hydrostatic model, but recover an inner bump in the dynamic model.
@@ 6.1 Physical implications of the bumps/shadows
Figure 11: Top: Radial profile of the flaring angle $\varphi$ at the irradiation absorption surface showing decreased incidence of stellar rays in shadowed regions. Middle: Radial profiles of the surface aspect ratio $h$ at the $\tau_{\star}=1$ surface. We see that the profiles deviate from a smooth power-law. Bottom: The midplane radial pressure gradient shows deviations in the pressure profile. We note the flaring angle $\varphi$ and the midplane pressure gradient $\text{d log}\penalty 10000\ P_{\text{mid}}/{\text{d log}\penalty 10000\ R}$ are smoothed with a rolling average over 20 cells to eliminate noise.
@@
To understand if the staircase structure we see in our radiation hydrodynamic simulations can produce observable features, in addition to the relative midplane temperature profiles and the temperature gradients plotted in Figs. 6 and 9, we provide further analysis here. The first two panels of Fig. 11 show the radial profiles of the flaring angle $\varphi$ and the height of the irradiation surface normalized to the local radius $z_{s}/R\text{ where }z_{s}=z\penalty 10000\ (\tau_{\star}=1)$. The flaring angle (i.e., the angle at which stellar rays hit the surface of the disk) is defined similar to the classical approach of 1997ChiangGoldreich and 2000Dullemond, where $\varphi=\sin\left[\nicefrac{{\text{d}z_{s}}}{{\text{d}R}}-\arctan(\nicefrac{{z_{s}}}{{R}})\right]$. We use the above mentioned quantities to trace whether the bumps/shadows produced in our models can affect the optical/NIR intensities producing observable signatures. We see that $\varphi$ decreases downstream of the bumps, blocking stellar rays to efficiently reach the shadows. This is also reflected by the non-uniform profiles of $z_{s}/R$, deviating from a smooth power-law profile. This provides evidence that the optical/NIR intensities are indeed reduced in the shadowed regions, but to also illustrate this further we provide synthetic scattered light images in the H Band (VLT/SPHERE) with RADMC-3D (2012Dullemond) in Appendix. B showing that our models produce bright rings and subsequent dark shadows. We also plot the midplane pressure gradient for our models in the bottom panel of Fig. 11 to understand the implications of the shadows/bumps in our models in the mm continuum. We do not see the inversion of the pressure gradient (i.e., dust traps, 1972Whipple; 2004Paardekooper) for any of our models, but we see deviations in the pressure gradient at the shadow locations. This implies that while the staircase structure due to the starlight-shadowing mechanism is not strong enough to trap dust, they could act as dust ”traffic jams” that could slow down dust grain drift, and also act as seeds for secondary dust accumulation mechanisms (2021JiangOrmel). We highlight that the self-consistent modeling of dust–gas dynamics coupled with radiation transfer is required for more accurate constraints on observable signatures of shadows, which is why providing state-of-the-art synthetic observations both in NIR and the mm continuum is not the focus of this work.
@@ Acknowledgements.
PS thanks Satoshi Okuzumi and Thomas Pfeil for helpful comments. PS is grateful for the many discussions on steady states and equilibria with Siva Athreya. PS acknowledges the support of the German Science Foundation (DFG) through grant number 495235860 and is a Fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). MF acknowledges support from the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 757957). AZ and TB acknowledge funding from the European Union under the European Union’s Horizon Europe Research and Innovation Programme 101124282 (EARLYBIRD). Views and opinions expressed are those of the authors only. All plots in this paper were made with the Python library matplotlib (2007Hunter). Typesetting was expedited with the use of GitHub Copilot and ChatGPT, but without the use of any AI-generated text.
@@ Appendix A Comparing FLD to M1 moment transfer
Both the M1 and the FLD methods overestimate the temperature at the midplane, which is expected to be lower using Monte Carlo methods (2002Dullemond; 2025Krieger; 2025Fuksman). This is also clear in in Fig. 13 by the comparison of the FLD model (blue dotted line) to the corresponding RADMC-3D calculation (green dotted line). 2024Kutra argued that this casts doubt on the validity of the results in 2022FuksmanKlahr, suggesting that this effect may preclude the formation of a temperature staircase. However, despite this overestimation, both FLD and M1 models reproduce shadowed structures in hydrostatic simulations, and we find a parameter space where shadows persist in hydrodynamic models with FLD. Thus, we find the argument presented in 2024Kutra to lack strong physical motivation. Instead, the presence or absence of a staircase is sensitive to the assumed dust content, the production of initial bumps, and the correct treatment of diffusive cooling timescales resulting from the balance of stellar irradiation and thermal cooling.
@@ Appendix C Thermal relaxation timescale
Figure 15: Left: Radial profile of the cooling timescale $\beta$ for the disk. The regions of the disk where we find starlight-driven shadows lie in the intermediate cooling timescale regime (1301-30\,au).Right: Vertical profile of the cooling timescale $\beta$ at $R=1$ au. We see that for most of the disk (except for the atmosphere, and outermost regions), optically thick cooling is more dominant.

### v2

@@ Abstract
Context. Protoplanetary disks observed in millimeter continuum and scattered light show a variety of substructures. While embedded planets are a common explanation, during the early planet formation phase, various physical processes in the disk could also trigger such features. One such possibility that has been previously theorized for passive disks is the thermal wave instability or the stellar irradiation instability—the flared disk may become unstable as directly illuminated regions puff up and cast shadows behind them. This would manifest as bright and dark rings, and a staircase-like structure in the disk optical surface.
@@
Conclusions. We conclude that we can recover thermally induced flared-staircase structures in radiation hydrodynamic simulations of irradiated protoplanetary disks using the flux-limited diffusion method. We find that the shadowing effect is sensitive to the dust content in the disk. We highlight the importance of modeling dust dynamics consistently to explain starlight-driven shadows.
@@ 1 Introduction
Disk substructures are ubiquitous in modern observations of planet-forming disks across different wavelengths. These include millimeter thermal emission from ground-based instruments like ALMA (2018Andrews), scattered light observations with VLT-SPHERE (2018Avenhaus; 2023Benisty), and more recently, studies with JWST (2025Tazaki). Planets in the disk could carve out gaps (2018Zhang) that explain substructure, of which the planetary gaps in the PDS 70 and WISPIT 2 systems are direct observational constraints (2018Keppler; 2019Haffert; 2025VanCap; 2025Close). Many physical processes in the disk could also trigger substructures (2023Bae), such as zonal flows due to the magneto-rotational instability (MRI, 2011Uribe; 2015Flock), spirals launched by gravitational instability (GI, 2016Kratter), gaps and vortices resulting from the vertical shear instability (VSI, 2020Flock), and spirals and rings as a consequence of misaligned inner disks (2024Zhang; 2025bZiampras) among others. Mechanisms such as these which could invert the local pressure gradient, radially pile up dust in the disk, acting as dust traps (1972Whipple; 2004Paardekooper; 2026Ormel).
@@
The mechanism behind this starlight-driven shadowing is outlined as follows (see also Fig. 1 in 2021Ueda, for an illustration): in vertical hydrostatic equilibrium, the balance between gravity and pressure— set by stellar irradiation heating– yields a passive (irradiated) disk with flared geometry. Thus, outer regions absorb substantially more stellar light than in an internally-heated, conical disk. For optically thick disks, scale height perturbations could trigger thermal waves that travel inwards, puffing up regions of the disk and casting shadows behind them. This shows up as bright and dark bands or shadows in temperature, or staircases in the optical surface. One can refer to them as ”unstable” shadows to distinguish them from the ”stable” shadow cast by the hot inner rim in transition disks (2001Dullemond) or due to an opacity transition at the snow line (2015Bailli).
@@
There have been some efforts in modeling the thermal wave instability with 2.5D radiation hydrodynamic simulations in recent literature. 2022Pavlyuchenkov2D used a diffusion solver for radiation transport and 2022FuksmanKlahr used M1 moment transfer coupled with frequency-dependent irradiation. While these works reported inwardly propagating thermal waves and shadows in the hydrostatic models, both studies concluded that the instability is damped in dynamical simulations. 2022FuksmanKlahr attributed this to the inefficiency in transporting thermal perturbations from the atmosphere to the midplane in the disk, as radial advection is more dominant. 2024Kutra used a simpler ”radiative forcing” method which enforced the vertically isothermal approximation and found staircase structures that did not propagate inwards but settled in the disk for several orbital timescales. Recently, 2025Mori also reported thermal wave instability-like behavior in their global non-ideal magneto-hydrodynamic models, particularly in the one including the Hall effect and a magnetic field with anti-aligned polarity.
@@
Figure 1: Schematic to illustrate the different terms: ”bump”, ”shadow” and ”staircase” defined in the introduction. The star is denoted by the orange blob, and the stellar rays hit the irradiation absorption surface $\tau_{\star}=1$, at height $z_{s}=z\penalty 10000\ (\tau_{\star}=1)$ from the midplane, and a flaring angle $\varphi$.
For clarity throughout this article, we adopt the following definitions for the words ”staircase”, ”bump”, and ”shadow”: a staircase is a step-like modulation in the irradiation absorption surface as a function of radius, consisting of a heated, puffed-up front (a bump) and a colder backside (a shadow). In this context, we note that a bump refers to the locally elevated, irradiated region, rather than the more traditional temperature/pressure gradient inversion commonly associated with the term. A shadow, in turn, corresponds to the dip in the irradiation flux downstream, which we trace through the midplane temperature decrease of at least 5% relative to the background power-law, this is clearly illustrated by the sketch in Fig. 1. In Sect. 2 we outline the physics used in our model, the model parameters, and the numerical methods used. In Sects. 35 we show the results from our hydrostatic, dynamical, and dust models. We then discuss the implications of our results in the context of previous work in Sect. 6 and summarize our findings in Sect. 7.
@@ 2.1 Radiation hydrodynamic equations
The gravitational potential of the star is given by $\Phi_{\star}=-\text{G}M_{\star}/r$ where G is the gravitational constant, $M_{\star}$ is the mass of the central star, and $r=\sqrt{R^{2}+z^{2}}$ is the spherical radius distance to the star. The total energy is a function of kinetic energy and the specific internal energy $e$ with $E_{\text{tot}}=\rho e+\frac{1}{2}\rho\bm{u}^{2}$. The gas pressure is denoted by $P$, and $\bm{\mathrm{I}}$ is the unit matrix. The temperature-dependent Planck and Rosseland mean opacities are $\kappa_{\text{P}}$ and $\kappa_{\text{R}}$ respectively, $a_{\text{R}}$ is the radiation constant, and $c$ is the speed of light. The set of equations is complete with a closure relation, that is, the ideal gas equation $P=(\gamma-1)\rho e$ with the adiabatic index chosen as $\gamma=1.42$ (1978Decampli) and the mean molecular weight $\mu=2.35$. The gas temperature $T$ is given by $e=\rho c_{\text{V}}T$, with $c_{\text{V}}=R_{\text{gas}}/\mu(\gamma-1)$ being the specific heat capacity at constant volume, and $R_{\text{gas}}$ the gas constant. The isothermal sound speed is defined as $c_{\mathrm{s}}=\sqrt{P/\rho}$ and the gas pressure scale height is given by $H=c_{\mathrm{s}}/\Omega_{\mathrm{K}}$ where $\Omega_{\mathrm{K}}=\sqrt{GM_{\star}/R^{3}}$ is the Keplerian angular velocity.
@@ 2.3 Model parameters
Table 1: Parameters for the different hydrostatic and dynamic runs. [^1] Each model is denoted by M followed by a label indicating the parameter altered relative to the fiducial run MFID.
@@
ML1∗*∗**ML1 is an additional model we set up that matches 2022FuksmanKlahr with a stellar parameters corresponding to $L_{\star}=1\,\mathrm{L}_{\odot}$, that we only discuss during the comparison in Appendix A. $0.01$ 11 $4000$ $2.086$ $200$
@@ 4 Results II: Changing the dust content
In this section, we outline our results from our hydrostatic and dynamical models for the low dust-to-gas ratio case of $\epsilon_{\text{dg}}=0.001$. As outlined above, these values are more typical to the dust content observed in later stages (i.e., Class II protoplanetary disks 2021Rilinger; 2024Birnstiel). Figure 7 shows the temperature and the irradiation source term after 30 iterations for the hydrostatic model (top) and the hydrodynamic model (bottom) after $t=500$ orbits. We do not see any temperature bumps/shadows forming in the hydrostatic model, but recover an inner bump in the dynamic model.
@@ 6.1 Physical implications of the bumps/shadows
Figure 11: Top: Radial profile of the luminosity factor $\varphi$ at the irradiation absorption surface showing decreased incidence of stellar rays in shadowed regions. Middle: Radial profiles of the surface aspect ratio $h$ at the $\tau_{\star}=1$ surface. We see that the profiles deviate from a smooth power-law. Bottom: The midplane radial pressure gradient shows deviations in the pressure profile. We note the flaring angle $\varphi$ and the midplane pressure gradient $\text{d log}\penalty 10000\ P_{\text{mid}}/{\text{d log}\penalty 10000\ R}$ are smoothed with a rolling average over 20 cells to eliminate noise.
To understand if the staircase structure we see in our radiation hydrodynamic simulations can produce observable features, in addition to the relative midplane temperature profiles and the temperature gradients plotted in Figs. 6 and 9, we provide further analysis here. The first two panels of Fig. 11 show the radial profiles of the luminosity factor $\varphi=\text{sin}\,\alpha$, where $\alpha$ is the flaring (grazing) angle and the height of the irradiation surface normalized to the local radius $z_{s}/R\text{ where }z_{s}=z\penalty 10000\ (\tau_{\star}=1)$. The flaring angle is the angle at which stellar rays hit the surface of the disk defined similar to the classical approach of 1997ChiangGoldreich and 2000Dullemond, where $\alpha=\arctan\left(\nicefrac{{\text{d}z_{s}}}{{\text{d}R}}\right)-\arctan\,(\nicefrac{{z_{s}}}{{R}})$ and the irradiation factor $\varphi$ is the fraction of luminosity intercepted by the disk surface, essentially approximating Eq. (3) as
$F_{\star}\approx\phi\frac{L_{\star}}{4\pi r^{2}}.$ (10)
@@
For small angles (i.e., razor-thin disks), $\varphi\simeq\alpha$. We use the above mentioned quantities to trace whether the bumps/shadows produced in our models can affect the optical/NIR intensities producing observable signatures. We see that $\varphi$ decreases downstream of the bumps, blocking stellar rays to efficiently reach the shadows. This is also reflected by the non-uniform profiles of $z_{s}/R$, deviating from a smooth power-law profile. This provides evidence that the optical/NIR intensities are indeed reduced in the shadowed regions, but to also illustrate this further we provide synthetic scattered light images in the H Band (VLT/SPHERE) with RADMC-3D (2012Dullemond) in Appendix. B showing that our models produce bright rings and subsequent dark shadows. We also plot the midplane pressure gradient for our models in the bottom panel of Fig. 11 to understand the implications of the shadows/bumps in our models in the mm continuum. We do not see the inversion of the pressure gradient (i.e., dust traps, 1972Whipple; 2004Paardekooper) for any of our models, but we see deviations in the pressure gradient at the shadow locations. This implies that while the staircase structure due to the starlight-shadowing mechanism is not strong enough to trap dust, they could act as dust ”traffic jams” that could slow down dust grain drift, and also act as seeds for secondary dust accumulation mechanisms (2021JiangOrmel). We highlight that the self-consistent modeling of dust–gas dynamics coupled with radiation transfer is required for more accurate constraints on observable signatures of shadows, which is why providing state-of-the-art synthetic observations both in NIR and the mm continuum is not the focus of this work.
@@ Acknowledgements.
The authors thank the anonymous referee for a thorough report, corrections and comments that improved this manuscript. PS thanks Satoshi Okuzumi and Thomas Pfeil for helpful comments. PS is grateful for the many discussions on steady states and equilibria with Siva Athreya. PS acknowledges the support of the German Science Foundation (DFG) through grant number 495235860 and is a Fellow of the International Max Planck Research School for Astronomy and Cosmic Physics at the University of Heidelberg (IMPRS-HD). MF acknowledges support from the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 757957). AZ and TB acknowledge funding from the European Union under the European Union’s Horizon Europe Research and Innovation Programme 101124282 (EARLYBIRD). Views and opinions expressed are those of the authors only. All plots in this paper were made with the Python library matplotlib (2007Hunter). Typesetting was expedited with the use of GitHub Copilot and ChatGPT, but without the use of any AI-generated text.
@@ Appendix A Comparing FLD to M1 moment transfer
Both the M1 and the FLD methods overestimate the temperature at the midplane, which is expected to be lower using Monte Carlo methods (2002Dullemond; 2025Krieger; 2025Fuksman; 2025Pavlyuchenkov). This is also clear in in Fig. 13 by the comparison of the FLD model (blue dotted line) to the corresponding RADMC-3D calculation (green dotted line). 2024Kutra argued that this casts doubt on the validity of the results in 2022FuksmanKlahr, suggesting that this effect may preclude the formation of a temperature staircase. However, despite this overestimation, both FLD and M1 models reproduce shadowed structures in hydrostatic simulations, and we find a parameter space where shadows persist in hydrodynamic models with FLD. Thus, we find the argument presented in 2024Kutra to lack strong physical motivation. Instead, the presence or absence of a staircase is sensitive to the assumed dust content, the production of initial bumps, and the correct treatment of diffusive cooling timescales resulting from the balance of stellar irradiation and thermal cooling.
@@ Appendix C Thermal relaxation timescale
Figure 15: Top: Radial profile of the cooling timescale $\beta$ for the disk. The regions of the disk where we find starlight-driven shadows lie in the intermediate cooling timescale regime (1301-30\,au). Bottom: Vertical profile of the cooling timescale $\beta$ at $R=1$ au. We see that for most of the disk (except for the atmosphere, and outermost regions), optically thick cooling is more dominant.